OpenStudy (anonymous):

e^(ln x) = 2 solve the equation for x

OpenStudy (anonymous):

similar to previous problem take logarithms and try on your own! If any difficulties post here!

OpenStudy (anonymous):

e^(ln x ) = 2 => x = 2

OpenStudy (anonymous):

hey thanks i did it right then!

OpenStudy (anonymous):

y = e ^(ln x) MEANS ln x = ln y by definition.

OpenStudy (anonymous):

Remember e^ln x = x, and 10^logx = x

OpenStudy (anonymous):

yeah ok

OpenStudy (anonymous):

@pepsi e^x = 4, take ln of both sides, x = ln 4.